OpenStudy (anonymous):

x3 + x3 + x3 + x3 = ?

OpenStudy (♪chibiterasu):

I think you mean x^3?

OpenStudy (anonymous):

is it \[x^3 +x^3 +x^3 = ?\] @Alipooh then answer will be 4 x3

OpenStudy (anonymous):

yea .

OpenStudy (anonymous):

its boobo from my side... 3x^3

OpenStudy (anonymous):

no its meant to be x^3

OpenStudy (anonymous):

SO 4 of them

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^12

OpenStudy (♪chibiterasu):

@lalala101 We're adding like terms actually.

OpenStudy (♪chibiterasu):

We would only do that if it were multiplication :P

OpenStudy (anonymous):

thik it as simple as this,,, 4 of x^3 ie 4x^3

OpenStudy (jdoe0001):

\(\bf x^3+x^3+x^3+x^3\qquad say\qquad x^3=a\qquad thus\\ \quad \\ a+a+a+a=\square ?\)

OpenStudy (anonymous):

i'm confused LOL hmm.....

OpenStudy (anonymous):

x^3?

OpenStudy (anonymous):

i got it. just needed conformation :D

OpenStudy (anonymous):

wlcome to OS @Alipooh enjy your stay here

OpenStudy (anonymous):

lol

OpenStudy (♪chibiterasu):

I haven't been able to welcome anyone during the Orange War because I can't see Smart Scores :P